A total of $11,000 is invested in two funds paying 7% and 8% simple interest. If the yearly interest is $865, find the amount invested at each rate.
Answer by mananth(12270)

A total of $11,000 is invested in two funds paying 7% and 8% simple interest. If the yearly interest is $865, find the amount invested at each rate.
...
7% ------ $x
and 8% --------$11000-x
simple interest= $865
..
0.07x+0.08(11000-x)=865
multiply by100
7x+88000-8x=86500
-x=-1500
x= 1500 @ 7%
$9500 @8%
m.ananth@hotmail.ca